# Velocity effects slightly mitigating the quantumness degradation of an Unruh-DeWitt detector

### v1

@@ IV.4 Complementarity relation
Besides, this duality is typically demonstrated in interference experiments menzel2012wave ; wang2021molecular ; yoon2021quantitative ; chen2022experimental . Substituting Eqs. (52) and (58) into (59), we concluded
@@
IV.5 Numerical results
Now, in this section we present the numerical results of the effects of non-relativistic velocity ($w\ll 1$) on the complementarity relation [Eq. (60)] obtained through interferometric visibility [Eq. (52)] and which-path distinguishability [Eq. (58)]. Through this analysis, it is possible to understand the wave-particle duality behavior of the system. In Fig. 4 we plot $C_{w}$ as a function of the $\overline{a}$ for distinct values of the $w$. Note that as we increase the acceleration the complementarity relation decays, this occurs due to the effects caused by Unruh radiation that degrade the wave-particle information of systems at high accelerations Pedro2024robustness ; pedro2025mitigating ; barros2025information . On the other hand, when we take increasing values of $w$ we see that the degradation of this information is mitigated slightly.
@@ VI Summary and conclusion
Evidently, in the expressions for the non-relativistic velocity regime, the presence of terms proportional to $w^{2}$ was observed for all quantum systems studied, which serves as a witness to non-relativistic velocity effects. On the other hand, for the ultra-relativistic velocity regime we observe that the transition rates are proportional to $w^{-4}$ and this causes the Unruh effect to be suppressed, and consequently, we do not see any effects in the accelerated quantum systems. It should be emphasized that an UDW detector following such a trajectory, in the ultrarelativistic limit, exhibits the same effect as that of a system interacting with its environment through the presence of boundaries under certain circumstances yang2018quantum ; Jin2015Electromagnetic ; liu2016protecting ; liu2016inhibiting . It is known that the total suppression of the detector response depends on many conditions, therefore, our findings depend on the model and do not have a general character.
Furthermore, the numerical results show that for all studied quantum systems, the information degradation decreases as we increase the parameter $w$, implying that the four-velocity component under very sensitive conditions serves as a protective agent for the information of these systems. This interesting behaviors happen because of the composite effect of both velocity and acceleration abdolrahimi2014velocity ; Liu2021relativistic ; xie2025weak , and in this work we witness these effects for other systems. More specifically, our findings showed that both for an accelerated single-qubit and for interferometric circuits, velocity effects cause mitigation in the degradation of the quantumness of these quantum systems in the high acceleration regime.
@@ Acknowledgements.
P.H.M. Barros acknowledges the Brazilian funding agency CAPES for financial support through grant No. 88887.674765/2022-00 (Doctoral Fellowship – CAPES); S.M. Wu was supported by the National Natural Science Foundation of China (Grant Nos. 12205133) and the Special Fund for Basic Scientific Research of Provincial Universities in Liaoning under grant No. LS2024Q002. C.A.S. Almeida is supported by grant number 309553/2021-0 (CNPq/PQ) and by Project UNI-00210-00230.01.00/23 (FUNCAP).

### v2

@@ 4.4 Complementarity relation
Besides, this duality is typically demonstrated in interference experiments Menzel et al. (2012); Wang and Yang (2021); Yoon and Cho (2021); Chen et al. (2022). Substituting Eqs. (50) and (56) into (57), we concluded
@@
4.5 Numerical results
Now, in this section we present the numerical results of the effects of non-relativistic velocity ($w\ll 1$) on the complementarity relation [Eq. (58)] obtained through interferometric visibility [Eq. (50)] and which-path distinguishability [Eq. (56)]. Through this analysis, it is possible to understand the wave-particle duality behavior of the system. In Fig. 7 we plot $C_{w}$ as a function of the $\overline{a}$ for distinct values of the $w$. Note that as we increase the acceleration the complementarity relation decays, this occurs due to the effects caused by Unruh radiation that degrade the wave-particle information of systems at high accelerations Barros et al. (2024, 2025b, 2025a). On the other hand, when we take increasing values of $w$ we see that the degradation of this information is mitigated slightly.
@@ 6 Summary and conclusion
Evidently, in the expressions for the non-relativistic velocity regime, the presence of terms proportional to $w^{2}$ was observed for all quantum systems studied, which serves as a witness to non-relativistic velocity effects. On the other hand, for the ultra-relativistic velocity regime we observe that the transition rates are proportional to $w^{-4}$ and this causes the Unruh effect to be suppressed, and consequently, we do not see any effects in the accelerated quantum systems. It should be emphasized that an UDW detector following such a trajectory, in the ultrarelativistic limit, exhibits the same effect as that of a system interacting with its environment through the presence of boundaries under certain circumstances Yang et al. (2018); Jin and Yu (2015); Liu et al. (2016b, a). It is known that the total suppression of the detector response depends on many conditions, therefore, our findings depend on the model and do not have a general character.
Furthermore, the numerical results show that for all studied quantum systems, the information degradation decreases as we increase the parameter $w$, implying that the four-velocity component under very sensitive conditions serves as a protective agent for the information of these systems. This interesting behaviors happen because of the composite effect of both velocity and acceleration Abdolrahimi (2014); Liu et al. (2021); Hao et al. (2023); Mukherjee et al. (2024); Xie et al. (2025), and in this work we witness these effects for other systems. More specifically, our findings showed that both for an accelerated single-qubit and for interferometric circuits, velocity effects cause mitigation in the degradation of the quantumness of these quantum systems in the high acceleration regime.
@@ Acknowledgements.
P.H.M. Barros acknowledges the Brazilian funding agency CAPES for financial support through grant No. 88887.674765/2022-00 (Doctoral Fellowship – CAPES); S.M. Wu was supported by the National Natural Science Foundation of China (Grant Nos. 12205133) and the Special Fund for Basic Scientific Research of Provincial Universities in Liaoning under grant No. LS2024Q002. C.A.S. Almeida is supported by grant number 309553/2021-0 (CNPq/PQ) and by Project UNI-00210-00230.01.00/23 (FUNCAP). The authors thank the Referee for helpful comments.
Appendix A Derivation of Wightman function expansions
In this appendix section we explicitly show the derivations of the asymptotic expansions of the Wightman function given by Eq. (12), presenting the two expansions, namely: for very small $w$ and for very large $w$.
A.1 Non-relativistic regime
Considering that the component of the four-velocity $w$ is very small, that is, $w\ll 1$, then the non-relativistic regime is obtained. Therefore, for these conditions using the binomial expansion, we have that Eq. (11) can be written as $\alpha\approx a[1-\frac{w^{2}}{2}+\mathcal{O}(w^{4})]$, and substituting this expression into the Wightman function Eq. (12), we obtain
$\mathcal{W}_{w}^{+}(\Delta\tau)\approx-\frac{a^{2}(1-\frac{w^{2}}{2})^{4}}{16\pi^{2}}\Bigg\{\sinh^{2}{\Bigg[\frac{a}{2}\left(1-\frac{w^{2}}{2}\right)\left(\Delta\tau-2i\epsilon\right)\Bigg]}-\frac{w^{2}(a\Delta\tau)^{2}(1-\frac{w^{2}}{2})^{4}}{4}\Bigg\}^{-1}.$
Using the expansion given by $\left(1-\frac{w^{2}}{2}\right)^{4}\approx 1-2w^{2}+\mathcal{O}(w^{4})$, and disregarding the terms $\mathcal{O}(w^{4})$, we have
$\mathcal{W}_{w}^{+}(\Delta\tau)\approx-\frac{a^{2}(1-2w^{2})}{16\pi^{2}}\Bigg\{\sinh^{2}{\Bigg[\frac{a}{2}\left(\Delta\tau-2i\epsilon\right)-\frac{a\,w^{2}}{4}\left(\Delta\tau-2i\epsilon\right)\Bigg]}-\frac{w^{2}(a\Delta\tau)^{2}}{4}\Bigg\}^{-1}.$
Now, to proceed, we can make the following definitions:
$A\equiv\frac{a}{2}\left(\Delta\tau-2i\epsilon\right),\qquad\text{and}\qquad B\equiv\frac{a\,w^{2}}{4}\left(\Delta\tau-2i\epsilon\right).$ (62)
Thus, we can analyze the following hyperbolic trigonometric relationship,
$\sinh{(A-B)}=\sinh{(A)}\cosh{(B)}-\cosh{(A)}\sinh{(B)},$ (63)
where $B$ is very small, since it is proportional to $w^{2}$, thus the following approximations are valid:
$\cosh{(B)}\approx 1+\frac{B^{2}}{2},\qquad\text{and}\qquad\sinh{(B)}\approx B.$ (64)
Thus, we have
$\sinh{(A-B)}$ $\approx$ $\sinh{(A)}\left(1+\frac{B^{2}}{2}\right)-B\cosh{(A)},$ (65)
$\approx$ $\sinh{(A)}+\frac{B^{2}}{2}\sinh{(A)}-B\cosh{(A)},$ (66)
where $B^{2}\sim\mathcal{O}(w^{4})$, and therefore, we obtain
$\sinh{(A-B)}$ $\approx$ $\sinh{(A)}\left(1+\frac{B^{2}}{2}\right)-B\cosh{(A)},$ (67)
$\approx$ $\sinh{(A)}-B\cosh{(A)}+\mathcal{O}(w^{4}).$ (68)
Now, we can square both sides of this expression, disregarding the terms $\mathcal{O}(w^{4})$, and thus we obtain:
$\sinh^{2}{(A-B)}\approx\sinh^{2}{(A)}-2B\sinh{(A)}\cosh{(A)}+\mathcal{O}(w^{4}),$ (69)
where using the hyperbolic trigonometric relation given by $\sinh{(2A)}=2\sinh{(A)}\cosh{(A)}$, we have
$\sinh^{2}{(A-B)}\approx\sinh^{2}{(A)}-B\sinh{(2A)}+\mathcal{O}(w^{4}).$ (70)
Now, substituting the relation given by Eq. (70) into Eq. (61), we have that the Wightman function is written as
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}(1-2w^{2})}{16\pi^{2}}\Bigg\{\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}-w^{2}\left(\frac{\Delta\tau}{4}-i\epsilon a\right)\sinh{\left(a\Delta\tau-2i\epsilon a\right)}+$ (71)
- $\frac{w^{2}(a\Delta\tau)^{2}}{4}\Bigg\}^{-1}+\mathcal{O}(w^{4}).$
To proceed with the mathematical manipulations in a clearer and simpler way, we can make the following definitions:
$D$ $\equiv$ $\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)},$ (72)
$E$ $\equiv$ $w^{2}\left(\frac{\Delta\tau}{4}-i\epsilon a\right)\sinh{\left(a\Delta\tau-2i\epsilon a\right)},$ (73)
$F$ $\equiv$ $\frac{w^{2}(a\Delta\tau)^{2}}{4},$ (74)
and with these definitions, our Wightman function now depends on the following term, $(D-E-F)^{-1}$, on which we can perform the following manipulations:
$\frac{1}{D-E-F}=\frac{1}{D-E-F}\left(\frac{D+E+F}{D+E+F}\right)=\frac{D+E+F}{D^{2}-(E+F)^{2}}.$ (75)
Note that the term $(B+C)^{2}$ is of the order of $\mathcal{O}(w^{4})$ and therefore we can disregard it, and thus we can continue with the following algebraic manipulations:
$\frac{1}{D-E-F}\approx\frac{D+E+F}{A^{2}}+\mathcal{O}(w^{4})\approx\frac{1}{D}+\frac{E+F}{D^{2}}+\mathcal{O}(w^{4}),$ (76)
and replacing the definitions given by Eqs. (72), (73), and (74), the Wightman function can be read as
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}(1-2w^{2})}{16\pi^{2}}\Bigg\{\frac{1}{\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}+\Bigg[w^{2}\left(\frac{\Delta\tau}{4}-i\epsilon a\right)\sinh{\left(a\Delta\tau-2i\epsilon a\right)}+$ (77)
$+$ $\frac{w^{2}(a\Delta\tau)^{2}}{4}\Bigg]\frac{1}{\sinh^{4}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}\Bigg\}+\mathcal{O}(w^{4}).$
Finally, we have the expansion of the Wightman function for $w\ll 1$, and it is written as
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}}{16\pi^{2}}\Bigg\{\frac{(1-2w^{2})}{\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}+\Bigg[\left(\frac{\Delta\tau}{4}-i\epsilon a\right)\sinh{\left(a\Delta\tau-2i\epsilon a\right)}+$ (78)
$+$ $\frac{(a\Delta\tau)^{2}}{4}\Bigg]\frac{w^{2}}{\sinh^{4}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}}\Bigg\}+\mathcal{O}(w^{4}),$
where in the last step we multiply the term $(1-2w^{2})$ across all terms within parentheses and disregard the term $\mathcal{O}(w^{4})$.
A.2 Ultra-relativistic regime
Now, in this section we derive in detail the expansion of the Wightman function for the ultra-relativistic regime, that is, when $w\gg 1$. Under these conditions, it is easy to see that for very large speeds the parameter $\alpha=\frac{a}{\sqrt{1+w^{2}}}$ can now be written as $\alpha\approx\frac{a}{w}$, and applying this to Eq. Eq. (12), we obtain
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}}{16\pi^{2}}\frac{1}{w^{4}}\Bigg[\sinh^{2}{\left(\frac{a\Delta\tau}{2w}-\frac{i\epsilon a}{w^{2}}\right)}-\frac{(a\Delta\tau)^{2}}{4w^{2}}\Bigg]^{-1},$ (79)
and to proceed, we can make the following definitions:
$G$ $\equiv$ $\sinh^{2}{\left(\frac{a\Delta\tau}{2w}-\frac{i\epsilon a}{w^{2}}\right)},$ (80)
$H$ $\equiv$ $\frac{(a\Delta\tau)^{2}}{4w^{2}}.$ (81)
Through these definitions, it is possible to note that the Wightman function now depends on the term given by
$\frac{1}{w^{4}}\frac{1}{G-H}=\frac{1}{w^{4}}\frac{1}{G-H}\left(\frac{G+H}{G+H}\right)=\frac{G+H}{w^{4}G^{2}-w^{4}H^{2}},$ (82)
where $w^{4}H^{2}$ is very small, since it is proportional to $w^{-8}$ and this term can be disregarded. Thus, we obtain
$\frac{1}{w^{4}}\frac{1}{G-H}\approx\frac{G+H}{w^{4}G^{2}}\approx\frac{1}{w^{4}G}+\frac{H}{w^{4}G^{2}},$ (83)
Note that the term $\frac{H}{w^{4}G^{2}}$ is proportional to $w^{-6}$, and we can disregard it as well, and therefore we have
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}}{16\pi^{2}}\frac{1}{w^{4}}\Bigg[\sinh^{2}{\left(\frac{a\Delta\tau}{2w}-\frac{i\epsilon a}{w^{2}}\right)}\Bigg]^{-1}+\mathcal{O}(w^{-6}),$ (84)
and this expression is the expansion of the Wightman function to the ultra-relativistic regime.